OpenStudy (anonymous):

What is the slope of a tangent line to f(x) = -2x^2 - 3x + 4 , at x = -2? Show work.

OpenStudy (anonymous):

slope =f'(x) at x=-2

OpenStudy (anonymous):

I have no idea how to do this, **EDIT TO QUESTION, or explain what steps I have to take and why(if applicable)

OpenStudy (anonymous):

take the first derivative of f(x) f'(x)=-4x - 3 at x=-2 f'(x)=8-3=5 so the slope is 5

OpenStudy (anonymous):

Thanks so much, I learned this on khanacademy , but wasn't sure if thats what I was suppose to do.